OpenStudy (anonymous):

It takes 3.2 × 10^3 J of heat to melt an amount of gold (△Hf = 6.30 × 10^4 J/kg). How much gold do you have? 0.051 g , 0.51 g , 19 g , or 51 g ?

OpenStudy (michele_laino):

please, for previous answer, I think inert gas, so not plasma

OpenStudy (michele_laino):

inert gas means a gas composed by neutral atoms

OpenStudy (anonymous):

ohh okay! so gas is the solution?

OpenStudy (michele_laino):

yes!

OpenStudy (anonymous):

ookay!! yay!! what about this one?

OpenStudy (michele_laino):

here we have to compute this: \[m = \frac{Q}{{\Delta {H_f}}}\] where m is the mass of melted gold

OpenStudy (anonymous):

ok! what do we plug in? :/

OpenStudy (michele_laino):

hint: \[m = \frac{Q}{{\Delta {H_f}}} = \frac{{3.2 \times {{10}^3}}}{{6.3 \times {{10}^4}}}\]

OpenStudy (anonymous):

okay! so we get 0.0507 so our solution is 0.051 g ?

OpenStudy (michele_laino):

no, it is 0.051 Kg

OpenStudy (anonymous):

ohh and so converted to g that becomes 0.51 g? :/

OpenStudy (anonymous):

i forget the conversoins :/

OpenStudy (michele_laino):

no, it is 0.051*1000=...grams

OpenStudy (anonymous):

ohh right so our solution is 51 g ? :O

OpenStudy (michele_laino):

yes!

OpenStudy (anonymous):

yay! thank you:)

OpenStudy (michele_laino):

:)